Tohban Report 2016-05-04
From RHESSI Wiki
|Start Date:||27 Apr 2016|
|End Date:||04 May 2016|
|Next Tohban:||Jim McTiernan|
|List all reports|
The activity has been low throughout the week, although there were 8 C-class flares. There are currently 7 active regions on disk and they are not showing any complex features. Also, 3 of them are expected to rotate off the disk within a few days.
How many GOES flares occurred?
Flares above B, C, M, X class were 45 8 0 0
RHESSI flare list contains 3 over the time range 26-Apr-16 03-May-16
No overlapping RHESSI flares to count.
Decimation was active/vigorous for the full time period. Detectors 1, 4 and 7 were turned on during the week. The SSR remained low, emptying at the end of daily pass sets. Spin rate is ~15rpm. Shadow length beginning to decrease, expected minimum on May 14.
Currently, the cold plates 1 and 2 are at 134.4 and 133.1. On 2016-04-27, the cold plates, 1 and 2 were at 134.3 and 132.9 K. Cryocooler power at 72 W. Acceleration is low at around 10mg. Cold tip 1 and 2 temperatures are 112 and 111 degrees.
High voltage turn on (Phase 1) started on April 27. See "Detector issues" below for more details. The power setting was kept at 95% at the time of the turn-on, but was raised to 100% on April 28. It was kept at 100% for the remainder of the week, slightly exceeding the 100% guideline as the temperatures continue to naturally increase due to the declining Sun fraction. Detector 1 is segmented and is being used for the attenuator logic. The threshold values are as follows, given in "% deadtime." A2 is shown for completeness, but is not currently used.
0.00 % Min A0 6.27 % Max A0 2.74 % Min A1 8.23 % Max A1 1.96 % Min A2 16.08 % Max A2 3.53 % Min A3 100.0 % Max A3
Adjustments to ADP were also made.
GAPS IN APP_ID = 102 (VC3-MONITOR RATES) WITH PACKET RATE LT 30
GAP START TIME GAP END TIME GAP (SEC) 2016-04-27T23:25:00.000 -- 2016-04-28T00:00:00.000 2100.0000 2016-04-28T00:00:00.000 -- 2016-04-28T00:50:00.000 3000.0000
High voltage turn on (Phase 1)
April 27: HV ramped up to 1985 V with no sign of segmentation.
April 28: Cryocooler power raised to 100% and the HV ramped to 3995 V. While D1 is segmented, D4 and D7 are not segmented and they continued to have high slow/fast valid rates.
May 03: HV on D4 & D7 ramped to the maximum voltage ~5000 V, but they were not segmented. Eventually, they were turned off, hoping to minimize additional heating.
So far, D1 is behaving well with the energy resolution of ~1.5 keV. As for threshold changes, on 2016 April 29, D4 Front fast and front-slow thresholds were raised to 0x38 and 0x30, respectively. Also, D7 Front fast and front-slow thresholds were raised to 0x50 and 0x70, respectively.
D3 and D5 are to be turned on after this Ops meeting.
|HLAT Decimation||Rear decimation weight 6, no front decimation|
|Night time data (fronts)||+/- 4 minutes|
|Night time data (rears)||+/- 4 minutes|
|Require extra passes?||No|
|Requirement for moving pointer?||No|
|Attenuator operation||Working as expected|
|Detector problems?||See notes above.|